OpenStudy (anonymous):

Complete the ordered pairs to solve: 4x + 3y = 8; (0,___); (___ ,0); (5,___)

OpenStudy (anonymous):

plug the values in 4(0)+3y=8 0+3y=8 divide by 3 so y would be 3/8 (0,3/8) since y=0 4x+3(0)=8 4x=8 x=2 (2,0) ____ 4(5)+3y=8 20+3y=8 3y=-12 y=-4